# Coarse-to-fine Q-Network with Action Sequence for Data-Efficient Robot Learning

### v5

@@ Abstract
Predicting a sequence of actions has been crucial in the success of recent behavior cloning algorithms in robotics. Can similar ideas improve reinforcement learning (RL)? We answer affirmatively by observing that incorporating action sequences when predicting ground-truth return-to-go leads to lower validation loss. Motivated by this, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a novel value-based RL algorithm that learns a critic network that outputs Q-values over a sequence of actions, i.e., explicitly training the value function to learn the consequence of executing action sequences. Our experiments show that CQN-AS outperforms several baselines on a variety of sparse-reward humanoid control and tabletop manipulation tasks from BiGym and RLBench.
Figure 1: Summary of results. Coarse-to-fine Q-Network with Action Sequence (CQN-AS) learns a critic network with action sequence. CQN-AS outperforms various RL and BC baselines such as CQN (Seo et al., 2024), DrQ-v2+ (Yarats et al., 2022), and ACT (Zhao et al., 2023) on 45 robotic tasks from BiGym (Chernyadev et al., 2024) and RLBench (James et al., 2020).
@@ 1 Introduction
Predicting action sequences from expert trajectories is a key idea in recent successful behavior cloning (BC; Pomerleau 1988) approaches in robotics. This has enabled policies to effectively approximate the noisy, multi-modal distribution of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Can this idea similarly be useful for reinforcement learning (RL)?
@@
Our initial finding is affirmative: we make an intriguing observation that using action sequences can enhance value learning. Specifically, with humanoid demonstrations from BiGym (Chernyadev et al., 2024), we train regression models that predict the ground-truth return-to-go, i.e., the sum of discounted future rewards from the timestep $t$, given the current observation and action. In Figure 2a, we find that using an action sequence $\mathbf{a}_{t:t+K}=\{\mathbf{a}_{t},...,\mathbf{a}_{t+K-1}\}$ as input results in lower validation losses than using a single-step action $\mathbf{a}_{t}$. We hypothesize this is because action sequences, which can correspond to behavioral primitives such as going straight, make it easier for the model to learn the long-term outcomes compared to evaluating the effect of individual single-step actions.
@@
Building on this observation, we train actor-critic algorithms (Haarnoja et al., 2018; Fujimoto et al., 2018) with action sequence on stand task from HumanoidBench (Sferrazza et al., 2024). Specifically, we train the actor to output action sequence and the critic to take action sequence as inputs instead of single-step actions. However, we find these algorithms with action sequences suffer from severe value overestimation (see Figure 2b) and completely fail to solve the task (see Figure 2c). This is because a wider action space makes the critic more vulnerable to function approximation error (Fujimoto et al., 2018) and the actor excessively maximizes value functions by exploiting this estimation error.
(a) Return-to-go prediction
(b) Value overestimation on stand
(c) Performance on stand
@@
Figure 2: Analyses. (a) We measure the improvement in the validation L1 loss of the return-to-go prediction model with different action sequence lengths. We find that using action sequence of length 50 results in the lower loss than using single-step action. (b) We find that SAC and TD3 with action sequences suffer from severe value overestimation in stand task from HumanoidBench (Sferrazza et al., 2024). (c) This leads to unstable training and failure to solve the task. These findings motivate us to design CQN-AS, a novel RL algorithm that incorporates action sequences for value learning while avoiding the value overestimation problem.
@@
This result motivates us to design our RL algorithm with action sequence upon a recent critic-only algorithm, i.e., Coarse-to-fine Q-Network (CQN; Seo et al. 2024), which solves continuous control tasks with discrete actions. Because there is no separate actor that may exploit value functions, i.e., CQN simply selects discrete actions with the highest Q-values, we find that training with action sequences is stable and thus avoids value overestimation problem. In particular, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), which learns a critic network that outputs Q-values over a sequence of actions (see Figure 3). By training the critic network to explicitly learn the consequence of taking a sequence of current and future actions, CQN-AS enables the RL agents to effectively learn useful value functions on challenging robotic tasks.
@@
Our experiments show that CQN-AS improves the performance of CQN on sparse-reward humanoid control tasks from BiGym benchmark (Chernyadev et al., 2024) that provides human-collected demonstrations and sparse-reward tabletop manipulation tasks from RLBench (James et al., 2020) that provide demonstrations generated via motion-planning. Considering that CQN-AS is a critic-only algorithm that selects actions with the highest Q-value without a separate actor network, these results highlight the benefit of using action sequences in value learning.
@@
  • We make an observation that shows using action sequences can be useful for RL by enhancing value learning. We also show that standard actor-critic algorithms (Haarnoja et al., 2018; Fujimoto et al., 2018) suffer from value overestimation when trained with action sequences.
@@
  • In a demo-driven RL setup that initializes training with expert demonstrations, we show that CQN-AS surpasses the performance of ACT (Zhao et al., 2023) – a BC algorithm that trains a Transformer (Vaswani et al., 2017) to predict action sequences.
@@ 2 Preliminaries
We formulate a robotic control problem as a partially observable Markov decision process (Kaelbling et al., 1998; Sutton & Barto, 2018). At time step $t$, an RL agent encounters an observation $\mathbf{o}_{t}$, executes an action $a_{t}$, receives a reward $r_{t+1}$, and encounters a new observation $\mathbf{o}_{t+1}$ from the environment. We aim to train a policy $\pi$ that maximizes the expected sum of rewards through RL while using as few online samples as possible, optionally with access to a modest amount of expert demonstrations.
@@
Figure 3: Coarse-to-Fine Q-Network with Action Sequence. CQN-AS extends Coarse-to-Fine Q-Network (CQN; Seo et al. 2024), a critic-only RL algorithm for continuous control using discretized actions. (a) CQN progressively zooms into the action space by discretizing it into $B$ bins and finding the bin with the highest Q-value to further discretize at the next level. Last level’s action sequence is used for controlling robots. CQN-AS generalizes this to action sequences by computing all $K$ actions in parallel. (b) We train a critic to predict Q-values over entire action sequences by extracting per-step features and aggregating them with a recurrent network before projection to Q-values.
@@
Coarse-to-fine Q-Network (CQN; Seo et al. 2024) is a critic-only RL algorithm that solves continuous control tasks with discrete actions. CQN trains an RL agent to learn to select coarse discrete actions in shallower levels with larger bin sizes, and then refine their choices by selecting finer-grained actions in deeper levels with smaller bin sizes. Specifically, CQN iterates the procedures of (i) discretizing the continuous action space into multiple bins and (ii) selecting the bin with the highest Q-value to further discretize. This reformulates the continuous control problem as a multi-level discrete control problem, allowing for the use of ideas from sample-efficient discrete RL algorithms (Mnih et al., 2015; Silver et al., 2017) for continuous control.
@@
[…]
where $\bar{\theta}$ are delayed parameters for a target network (Polyak & Juditsky, 1992) and $\pi^{l}$ is a policy that outputs the action $a_{t}^{l}$ at each level $l$ via the inference steps with our critic, i.e., $\pi^{l}(\mathbf{h}_{t})=a_{t}^{l}$. Specifically, to output actions at time step $t$, CQN first initializes constants $a_{t}^{\tt{low}}$ and $a_{t}^{\tt{high}}$ with $-1$ and $1$. Then the following steps are repeated for $l\in\{1,...,L\}$:
@@
    Step 1 (Discretization): Discretize an interval $[a_{t}^{\tt{low}},a_{t}^{\tt{high}}]$ into $B$ uniform intervals, and each of these intervals become an action space for $Q_{\theta}^{l}$.
@@
We then use the last level’s action $a_{t}^{L}$ as the action at time step $t$. For more details, including the inference procedure for computing Q-values, we refer readers to Appendix C.
@@ 3 Method
We present Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a value-based RL algorithm that learns a critic network that outputs Q-values for a sequence of actions $a_{t:t+K}=\{a_{t},...,a_{t+K-1}\}$ for a given observation $\mathbf{o}_{t}$. Our main motivation comes from one of the key ideas in recent BC approaches: predicting action sequences, which helps resolve ambiguity when approximating noisy distributions of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Similarly, by explicitly learning Q-values of a sequence of actions from the given state, our approach mitigates the challenge of learning Q-values with noisy trajectories. We provide the overview of CQN-AS in Figure 3.
@@ 3.1 Coarse-to-fine Critic with Action Sequence
[…]
where $N$ is a hyperparameter for $N$-step return and $\pi_{K}^{l}$ is an action sequence policy that outputs the action sequence $\mathbf{a}_{t:t+K}^{l}$ by following the similar inference procedure as in Section 2 (see Figure 3a). In practice, we compute Q-values for all sequence step $k\in\{1,...,K\}$ in parallel, which is possible as Q-values for future actions depend only on features $\mathbf{h}_{t}$ but not on previous actions.
@@
We note that any $N$-step return can be used in Equation 2 because the network can learn the long-term value of outputting action $a_{t+k}$ from bootstrapping. There is a trade-off: if one considers a short $N$-step return, it can cause a challenge as the setup becomes a delayed reward setup; but training with higher $N$-step return may introduce variance (Sutton & Barto, 2018). In our considered setups, we empirically find that using common values $N\in\{1,3\}$ works the best. We provide empirical analysis on the effect of $N$ in Figure 7c.
@@
Our critic network initially extracts features for each sequence step $k$ and aggregates features from multiple steps with a recurrent network (see Figure 3b). This architecture is helpful in cases where a single-step action is already high-dimensional so that concatenating them make inputs too high-dimensional. Specifically, let $\mathbf{e}_{k}$ denote an one-hot encoding for $k$. At each level $l$, we construct features for each sequence step $k$ as $\mathbf{h}^{l}_{t,k}=\left[\mathbf{h}_{t},a_{t+k-1}^{l-1},\mathbf{e}_{k}\right]$. We then encode each $\mathbf{h}_{t,k}^{l}$ with a shared MLP network and process them through GRU (Cho et al., 2014) to obtain $\mathbf{s}^{l}_{t,k}=f^{\tt{GRU}}_{\theta}(f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,1}^{l}),...,f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,k}^{l})$. We find that this design empirically performs better than directly giving actions as inputs to GRU. We then use a shared projection layer to map each $\mathbf{s}_{t,k}^{l}$ into Q-values at each sequence step $k$, i.e., $[Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t+k-1}^{l,b},a_{t:t+K}^{l-1})]_{b=1}^{B}=f_{\theta}^{\tt{proj}}(\mathbf{s}_{t,k}^{l})$.
@@ 3.2 Action Execution and Training Details
With the policy that outputs an action sequence $a_{t:t+K}$, one question is how to execute actions at time step $i\in\{t,...,t+K-1\}$. For this, we use temporal ensemble (Zhao et al., 2023) that computes $a_{t:t+K}$ every time step, saves it to a buffer, and executes a weighted average $\sum_{i}w_{i}\bar{a}_{t}^{i}/\sum w_{i}$ where $\bar{a}_{t}^{i}$ denotes an action for step $t$ computed at step $t-i$, $w_{i}=\exp(-m*i)$ denotes a weight that assigns higher value to more recent actions, with $m$ as a hyperparameter that adjusts the weighting magnitude. We find this scheme outperforms the alternative of computing $a_{t:t+K}$ every $K$ steps and executing each action for subsequent $K$ steps on most tasks we considered, except on several tasks that need reactive control.
@@
Figure 4: Examples of robotic tasks. We study CQN-AS on 25 humanoid control tasks from BiGym (Chernyadev et al., 2024) and 20 tabletop manipulation tasks from RLBench (James et al., 2020).

### v6

# Coarse-to-fine Q-Network with Action Sequence for Data-Efficient Reinforcement Learning

@@ Abstract
Predicting a sequence of actions has been crucial in the success of recent behavior cloning algorithms in robotics. Can similar ideas improve reinforcement learning (RL)? We answer affirmatively by observing that incorporating action sequences when predicting ground-truth return-to-go leads to lower validation loss. Motivated by this, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a novel value-based RL algorithm that learns a critic network that outputs Q-values over a sequence of actions, i.e., explicitly training the value function to learn the consequence of executing action sequences. Our experiments show that CQN-AS outperforms several baselines on a variety of sparse-reward humanoid control and tabletop manipulation tasks from BiGym and RLBench. Code is available at: https://younggyo.me/cqn-as/
Figure 1: Summary of results. Coarse-to-fine Q-Network with Action Sequence (CQN-AS) learns a critic network with action sequence. CQN-AS outperforms various RL and BC baselines such as CQN (seo2024continuous), DrQ-v2+ (yarats2022mastering), and ACT (zhao2023learning) on 45 robotic tasks from BiGym (chernyadev2024bigym) and RLBench (james2020rlbench).
@@ 1 Introduction
Predicting action sequences from expert trajectories is a key idea in recent successful behavior cloning (BC; pomerleau1988alvinn) approaches in robotics. This has enabled policies to effectively approximate the noisy, multi-modal distribution of expert demonstrations (zhao2023learning; chi2023diffusion). Can this idea similarly be useful for reinforcement learning (RL)?
@@
Our initial finding is affirmative: we make an intriguing observation that using action sequences can enhance value learning. Specifically, with humanoid demonstrations from BiGym (chernyadev2024bigym), we train regression models that predict the ground-truth return-to-go, i.e., the sum of discounted future rewards from the timestep $t$, given the current observation and action. In Figure˜2a, we find that using an action sequence $\mathbf{a}_{t:t+K}=\{\mathbf{a}_{t},...,\mathbf{a}_{t+K-1}\}$ as input results in lower validation losses than using a single-step action $\mathbf{a}_{t}$. We hypothesize this is because action sequences, which can correspond to behavioral primitives such as going straight, make it easier for the model to learn the long-term outcomes compared to evaluating the effect of individual single-step actions (see Appendix˜A for additional analysis based on a 2D Point-mass environment).
@@
Building on this observation, we train actor-critic algorithms (haarnoja2018soft; fujimoto2018addressing) with action sequence on stand task from HumanoidBench (sferrazza2024humanoidbench). Specifically, we train the actor to output action sequence and the critic to take action sequence as inputs instead of single-step actions. However, we find these algorithms with action sequences suffer from severe value overestimation (see Figure˜2b) and completely fail to solve the task. This is because a wider action space makes the critic more vulnerable to function approximation error (fujimoto2018addressing) and the actor excessively maximizes value functions by exploiting this estimation error. To further support this, in Figure˜2c, we report additional toy experiments where we introduce redundant no-op actions for training TD3 agents on Cheetah Run task (tassa2020dm_control). Here, we find that actor-critic algorithms are indeed vulnerable to value overestimation with high-dimensional action spaces.
Refer to caption (a) Return-to-go prediction Refer to caption (b) Value overestimation on stand TD3 CQN No-op added Return Value Return Value 0 140.7 16.25 236.6 18.8 54 7.18 -17.45 219.5 20.3 144 0.56 1E8 185.4 16.4 294 0.27 4E8 202.9 18.7 (c) Effect of no-op actions
@@
Figure 2: Analyses. (a) We measure the improvement in the validation L1 loss of the return-to-go prediction model with different action sequence lengths. We find that using action sequence of length 50 results in the lower loss than using single-step action. (b) We find that SAC and TD3 with action sequences suffer from severe value overestimation in stand task from HumanoidBench, which leads to random near-zero performance. (c) Actor-critic algorithms like TD3 become vulnerable to value overestimation when redundant no-op actions are added to the action space. In contrast, a critic-only algorithm that uses discrete actions, CQN, is robust with high-dimensional action spaces.
@@
This result motivates us to design our RL algorithm with action sequence upon a recent critic-only algorithm, i.e., Coarse-to-fine Q-Network (CQN; seo2024continuous), which solves continuous control tasks with discrete actions. Because there is no separate actor that may exploit value functions, i.e., CQN simply selects discrete actions with the highest Q-values, we find that training with action sequences is stable and thus avoids value overestimation problem (see Figure˜2c). In particular, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), which learns a critic network that outputs Q-values over a sequence of actions (see Figure˜3). By training the critic network to explicitly learn the consequence of taking a sequence of current and future actions, CQN-AS enables the RL agents to effectively learn useful value functions on challenging robotic tasks.
@@
Our experiments show that CQN-AS improves the performance of CQN on sparse-reward humanoid control tasks from BiGym benchmark (chernyadev2024bigym) that provides human-collected demonstrations and sparse-reward tabletop manipulation tasks from RLBench (james2020rlbench) that provide demonstrations generated via motion-planning. Considering that CQN-AS is a critic-only algorithm that selects actions with the highest Q-value without a separate actor network, these results highlight the benefit of using action sequences in value learning.
@@
  • We make an observation that shows using action sequences can be useful for RL by enhancing value learning. We also show that standard actor-critic algorithms (haarnoja2018soft; fujimoto2018addressing) suffer from value overestimation when trained with action sequences.
@@
  • In a demo-driven RL setup that initializes training with expert demonstrations, we show that CQN-AS surpasses the performance of ACT (zhao2023learning) – a BC algorithm that trains a Transformer (vaswani2017attention) to predict action sequences.
@@ 2 Preliminaries
We formulate a robotic control problem as a partially observable Markov decision process (kaelbling1998planning; sutton2018reinforcement). At time step $t$, an RL agent encounters an observation $\mathbf{o}_{t}$, executes an action $a_{t}$, receives a reward $r_{t+1}$, and encounters a new observation $\mathbf{o}_{t+1}$ from the environment. We aim to train a policy $\pi$ that maximizes the expected sum of rewards through RL while using as few online samples as possible, optionally with access to a modest amount of expert demonstrations.
@@
Figure 3: Coarse-to-Fine Q-Network with Action Sequence. CQN-AS extends Coarse-to-Fine Q-Network (CQN; seo2024continuous), a critic-only RL algorithm for continuous control using discretized actions. (a) CQN progressively zooms into the action space by discretizing it into $B$ bins and finding the bin with the highest Q-value to further discretize at the next level. Last level’s action sequence is used for controlling robots. CQN-AS generalizes this to action sequences by computing all $K$ actions in parallel. (b) We train a critic to predict Q-values over entire action sequences by extracting per-step features and aggregating them with a recurrent network before projection to Q-values.
@@
Coarse-to-fine Q-Network (CQN; seo2024continuous) is a critic-only RL algorithm that solves continuous control tasks with discrete actions. CQN trains an RL agent to learn to select coarse discrete actions in shallower levels with larger bin sizes, and then refine their choices by selecting finer-grained actions in deeper levels with smaller bin sizes. Specifically, CQN iterates the procedures of (i) discretizing the continuous action space into multiple bins and (ii) selecting the bin with the highest Q-value to further discretize. This reformulates the continuous control problem as a multi-level discrete control problem, allowing for the use of ideas from sample-efficient discrete RL algorithms (mnih2015human; silver2017mastering) for continuous control.
@@
$\begin{aligned} \mathcal{L}^{l}=\big(Q_{\theta}^{l}(\mathbf{h}_{t},a_{t}^{l},a_{t}^{l-1})-r_{t+1}-\gamma\max_{a^{\prime}}Q_{\bar{\theta}}^{l}(\mathbf{h}_{t+1},a^{\prime},\pi^{l-1}(\mathbf{h}_{t+1})\big),\end{aligned}$
where $\bar{\theta}$ are delayed parameters for a target network (polyak1992acceleration) and $\pi^{l}$ is a policy that outputs the action $a_{t}^{l}$ at each level $l$ via the inference steps with our critic, i.e., $\pi^{l}(\mathbf{h}_{t})=a_{t}^{l}$. Specifically, to output actions at time step $t$, CQN first initializes constants $a_{t}^{\tt{low}}$ and $a_{t}^{\tt{high}}$ with 1-1 and $1$. Then the following steps are repeated for $l\in\{1,...,L\}$:
@@
    Step 1 (Discretization): Discretize an interval $a_{t}^{\tt{low}},a_{t}^{\tt{high}}]$ into $B$ uniform intervals, and each of these intervals become an action space for $Q_{\theta}^{l}$.
@@
We then use the last level’s action $a_{t}^{L}$ as the action at time step $t$. For more details, including the inference procedure for computing Q-values, we refer readers to Appendix˜C.
@@ 3 Method
We present Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a value-based RL algorithm that learns a critic network that outputs Q-values for a sequence of actions $a_{t:t+K}=\{a_{t},...,a_{t+K-1}\}$ for a given observation $\mathbf{o}_{t}$. Our main motivation comes from one of the key ideas in recent BC approaches: predicting action sequences, which helps resolve ambiguity when approximating noisy distributions of expert demonstrations (zhao2023learning; chi2023diffusion). Similarly, by explicitly learning Q-values of a sequence of actions from the given state, our approach mitigates the challenge of learning Q-values with noisy trajectories. We provide the overview of CQN-AS in Figure˜3.
@@ 3.1 Coarse-to-fine Critic with Action Sequence
$\begin{aligned} \sum\nolimits_{k}\sum\nolimits_{l}\big(Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t+k-1}^{l},a_{t:t+K}^{l-1})-\sum\nolimits_{i=1}^{N}r_{t+i}-\gamma\max_{a^{\prime}}Q_{\bar{\theta}}^{l,k}(\mathbf{h}_{t+1},a^{\prime},\pi^{l-1}_{K}(\mathbf{h}_{t+1})\big)^{2},\end{aligned}$ (2)
where $N$ is a hyperparameter for $N$-step return and $\pi_{K}^{l}$ is an action sequence policy that outputs the action sequence $\mathbf{a}_{t:t+K}^{l}$ by following the similar inference procedure as in Section˜2 (see Figure˜3a). In practice, we compute Q-values for all sequence step $k\in\{1,...,K\}$ in parallel, which is possible as Q-values for future actions depend only on features $\mathbf{h}_{t}$ but not on previous actions.
@@
We note that any $N$-step return can be used in Equation˜2 because the network can learn the long-term value of outputting action $a_{t+k}$ from bootstrapping. There is a trade-off: if one considers a short $N$-step return, it can cause a challenge as the setup becomes a delayed reward setup; but training with higher $N$-step return may introduce variance (sutton2018reinforcement). In our considered setups, we empirically find that using common values $N\in\{1,4\}$ works the best. We provide empirical analysis on the effect of $N$ in Figure˜8c.
@@
Our critic network initially extracts features for each sequence step $k$ and aggregates features from multiple steps with a recurrent network (see Figure˜3b). This architecture is helpful in cases where a single-step action is already high-dimensional so that concatenating them make inputs too high-dimensional. Specifically, let $\mathbf{e}_{k}$ denote an one-hot encoding for $k$. At each level $l$, we construct features for each sequence step $k$ as $\mathbf{h}^{l}_{t,k}=\left[\mathbf{h}_{t},a_{t+k-1}^{l-1},\mathbf{e}_{k}\right]$. We then encode each $\mathbf{h}_{t,k}^{l}$ with a shared MLP network and process them through GRU (cho2014learning) to obtain $\mathbf{s}^{l}_{t,k}=f^{\tt{GRU}}_{\theta}(f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,1}^{l}),...,f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,k}^{l})$. We find that this design empirically performs better than directly giving actions as inputs to GRU. We then use a shared projection layer to map each $\mathbf{s}_{t,k}^{l}$ into Q-values at each sequence step $k$, i.e., $Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t+k-1}^{l,b},a_{t:t+K}^{l-1})]_{b=1}^{B}=f_{\theta}^{\tt{proj}}(\mathbf{s}_{t,k}^{l})$.
@@ 3.2 Action Execution and Training Details
With the policy that outputs an action sequence $a_{t:t+K}$, one question is how to execute actions at time step $i\in\{t,...,t+K-1\}$. For this, we use temporal ensemble (zhao2023learning) that computes $a_{t:t+K}$ every time step, saves it to a buffer, and executes a weighted average $\sum_{i}w_{i}\bar{a}_{t}^{i}/\sum w_{i}$ where $\bar{a}_{t}^{i}$ denotes an action for step $t$ computed at step tit-i, $w_{i}=\exp(-m*i)$ denotes a weight that assigns higher value to more recent actions, with $m$ as a hyperparameter that adjusts the weighting magnitude. We find this scheme outperforms the alternative of computing $a_{t:t+K}$ every $K$ steps and executing each action for subsequent $K$ steps on most tasks we considered, except on several tasks that need reactive control.
@@
Figure 4: Examples of robotic tasks. We study CQN-AS on 25 humanoid control tasks from BiGym (chernyadev2024bigym) and 20 tabletop manipulation tasks from RLBench (james2020rlbench).